5. Suppose a firm has the production function Q = K1/2 L1/4 M1/4 The wage rate w = 16 , rental rate r = 2 , and the price of the materials m = 1.

1) Suppose in the short run, K is fixed at K *. What’s the solution to the firm’s short run cost-minimization problem?

2) What is the solution to the firm’s long run cost minimization problem given that the firm wants to produce Q units of output?

3) Suppose that Q = 10, K* = 20. Compare the long run and short run demands.
Answer #1

Suppose a firm has the production function. Q: К% L4 м У . Now инв, in the short ren kis fined at k* да 41 24 - к+ + 2494 проNow, substituting M=16L in the Production function we gets att Q = k*%2/4 MY4 : a) Q = k* 1/2 2/4 (164)/4 =) Q **-1/2 = 22 =4 | Nov @ = к-чи, Now MPL = do 4 K/22 -3/4 m/4 MP = 20 = Ź K-²2 14 M | bet : , чи 4 - -) MPя 20 : +к 42 - 2 Now optimal inputNow set equation (ill equation (1) - к+-4 м 4 Тка. **и - — 5 кк, 6 - К% -/4 м _4 м -34 K42 ** m-34 =) ~ - 16 КИ. [-4 мм - 16Hnce ren demands are i the long K=2Q L = 20 M=20 3 suppose that k*=20, Q=10 Shord ren demands will be Ź 10 (20)/2 2 <= QK* 2